For the box of problem 1, Find how fast the volume increase as b increase


1
Expert's answer
2021-07-13T05:16:56-0400

Total Volume of box V = "b(b+1)(b+4)=b(b^2+5b+4)=b^3+5b^2+4b"


Differentiate V w.r.t. b-


"\dfrac{dV}{db}=3b^2+10b+4"


Hence Volume is increasing at "3b^2+10b+4"